OpenStudy (anonymous):

so what would this be x^-1 + y^-1 ______________ x^-3 +y^-3

myininaya (myininaya):

\[\frac{\frac{1}{x}+\frac{1}{y}}{\frac{1}{x^3}+\frac{1}{y^3}}\] remember \[a^3+b^3=(a+b)(a^2-ab+b^2)\]

myininaya (myininaya):

\[\frac{1}{x^3}+\frac{1}{y^3}=(\frac{1}{x}+\frac{1}{y})((\frac{1}{x})^2-\frac{1}{x} \cdot \frac{1}{y}+(\frac{1}{y})^2)\]

myininaya (myininaya):

do you see that the numerator and denominator have common factors?

myininaya (myininaya):

so now we have \[\frac{1}{\frac{1}{x^2}-\frac{1}{xy}+\frac{1}{y^2}}\]

OpenStudy (anonymous):

so multiply by the reciprocal?

myininaya (myininaya):

multiply top and bottom by \[x^2y^2\] to clear the fractions on bottom

OpenStudy (anonymous):

why x^2 y^2

myininaya (myininaya):

\[\frac{x^2y^2}{(x^2y^2)(\frac{1}{x^2}-\frac{1}{xy}+\frac{1}{y^2})}=\frac{x^2y^2}{y^2-xy+x^2}\]

myininaya (myininaya):

because we have x^2 and y^2 in the bottoms of the denominator

OpenStudy (anonymous):

lcd?

myininaya (myininaya):

the lcd of x^2,xy,y^2 is x^2y^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so is that the answer

OpenStudy (anonymous):

not the lcd i mean the fraction

OpenStudy (anonymous):

can it be simplified